Top answers


The probability of getting heads on a biased coin is 0.8. You flip the coin twice. What is the probability of getting one each of heads and tails?

Step 1: Calculate probability of getting tails by subtracting 1-0.8 = 0.2.Step 2: Calculate probability of getting heads then tails by multiplying 0.8 0.2 = 0.16.Step 3: Calculate probability of getting tail...
NR
Answered by Nadya R. • Maths tutor
11290 Views

How do you multiply two brackets with two terms in them? For example (2x-3)(x+4)

We use the FOIL method to do this, where FOIL stands for F irst, O utside, I nside, L ast. So in this example we multiply the first term of the first bracket with the first term of the second bracket ( 2x -3...
OA
Answered by Oluwatose A. • Maths tutor
4748 Views

Solve the following pair of simultaneous equations: 2y - x = 3, y + 4x = 4

There are 2 main methods for solving simultaneous equations: 1) by substitution, and 2) by elimination . Have a look at how each method works and pick the one that's best for you! In order to make my explana...
SP
Answered by Saranya P. • Maths tutor
4280 Views

Complete the square for the equation x^2 - 12x + 8 = 0

To complete the square, we will need to put the equation into the form (x - a) 2 - b + 8 = 0 , where a is half of the coefficient of x (12 in this case) and b is the value we need to subtract in order for th...
MJ
Answered by Mark J. • Maths tutor
4699 Views

Factorise (5x^2 + 7x + 2)

This expression is in the form of ax^2 + bx + c, so you need to find two numbers which multiply to get ac and add to make b. Step 1: Here a is not = 1, so you need to take out the highest common factor. In t...
DJ
Answered by Daisy J. • Maths tutor
11547 Views